OpenStudy (anonymous):

help please: write the following as a single log using logarithmic properties 6log5u + 4log5v

OpenStudy (anonymous):

would the correct answer be 24log(u)(v)

OpenStudy (anonymous):

where did you get the five from on the u?

OpenStudy (zzr0ck3r):

\(6\log_5(u)+4\log_5(v)=\log_5(u^6)+\log_5(v^4)=\log_5(u^6v^4)\)

OpenStudy (anonymous):

can you please explain one more problem to me please

OpenStudy (anonymous):

\[2\log_{b}x+1/2\log_{b}y+\log_{b}z \]

OpenStudy (anonymous):

alogb=logb^a here the base is e this is a logarithm property

OpenStudy (anonymous):

log a+log b = log(ab)

OpenStudy (anonymous):

so how would you do this equation \[2\log_{b}x+1/2\log_{b}y+\log_{b}z \]

OpenStudy (anonymous):

would it be like \[\log_{b}(x^6y^\frac{ 1 }{ 2 } z)\]

OpenStudy (anonymous):

@aayushi.somani

OpenStudy (anonymous):

how u got x^6

OpenStudy (anonymous):

sorry x^2

OpenStudy (anonymous):

and how u got y^1/2

OpenStudy (anonymous):

from the 1/2 thats in fron of log by the y

OpenStudy (anonymous):

good and how u got z

OpenStudy (anonymous):

@numnum